OpenStudy (anonymous):

Ugh...I hate these questions... :( Symbolically evaluate the given integrals. All work must be shown and proper notation must be used. (x+1)^2*(cos(x^3+3x^2+3x))dx Evaluated from 1 to 4

OpenStudy (anonymous):

you know how to use product rule ?

OpenStudy (asnaseer):

hint:\[\frac{d}{dx}(x^3+3x^2+3x)=3x^2+6x+3=3(x^2+2x+1)=3(x+1)^2\]and:\[\frac{d}{dx}(\sin(x))=\cos(x)\]